Im Working On An Online assignment on webassign.net. The Question is:
Use the functions given below to answer the following question.
f(x)=x^3, g(x)= x-3
Find' (f o f)(x) Answer by anantha(86) 
 sol:
(fof)(x)=f[f(x)]
=f[x^3] (in our problem f(x)=x^3 substitute this value in the place of f(x))
again substitute x^3 in place of x in x^3
f(x^3)=(x^3)^3
substitute x^3 in x
(fof)(x)=x^9 (i.e;(a^m)^n=a^mn)